OpenStudy (curry):

find an equivalent equation in retangular form for r-6sin=2cos

OpenStudy (ash2326):

\[ r- 6 \sin \theta= 2 \cos \theta\] Is this your question ???

OpenStudy (curry):

ye

OpenStudy (curry):

convert to rectangular fform

OpenStudy (ash2326):

You mean the form \[a+ib\]

OpenStudy (curry):

no like x^2 + y^2 form

OpenStudy (ash2326):

Ok, I got it \[r= 6 \sin \theta+ 2\cos \theta\] divide and multiply right hand side by by \(\sqrt{6^2+2^2}\) \[r=(6 \sin \theta+ 2\cos \theta)\times \frac{\sqrt {40}}{\sqrt {40}}\] \[r=\sqrt{40}(\frac 6{\sqrt 40} \sin \theta+ \frac 2{\sqrt {40}}\cos \theta)\] \[r=\sqrt {40} \sin ( \theta +x)\ \text{ where x=}\ \tan^{-1} \frac{2}{6}\]

OpenStudy (ash2326):

@curry Is this the form you need ?

OpenStudy (curry):

well ssomewhat in other word in the simplified form no r sshould be preent nor any sin or coss

OpenStudy (ash2326):

Is your \[r=x+iy\] ?

OpenStudy (curry):

here is my work r = 2cos+6sin r^2=2rcos+6rsin x^2+y^2=2x+6y

OpenStudy (curry):

i jussst dont get where to go from there

OpenStudy (ash2326):

You've done great \[x^2+y^2-2x-6y=0\] This is your rectangular equation

OpenStudy (ash2326):

@Curry I just realized that you've posted this in Biology. Please use maths for Maths' questions

OpenStudy (blues):

Hi Curry, please post math questions in the math group. Helps you get better answers (though Ash has given you a superb one) and helps us keep the site organised. Here is the link: http://openstudy.com/study?login#/groups/Mathematics. Thanks.